# Existence and large radial solutions for an elliptic system under finite new Keller-Osserman integral conditions

### v1

@@ Abstract
We investigate the existence and qualitative behavior of entire positive radial solutions to the semilinear elliptic system
@@
under finite Keller–Osserman-type integral conditions on the nonlinearities $f$ and $g$, and integrability constraints on the radial weights $p$ and $q$. The nonlinearities are assumed continuous on $0,\infty)$, differentiable on $(0,\infty)$, vanish at the origin, and are strictly positive elsewhere, with
$\int_{1}^{\infty}\frac{dt}{g(f(t))}<\infty,\qquad\int_{1}^{\infty}\frac{dt}{f(g(t))}<\infty.$
The weights satisfy $\int_{0}^{\infty}s\,p(s)\,ds<\infty$, $\int_{0}^{\infty}s\,q(s)\,ds<\infty$, and $\min(p,q)$ is not compactly supported. Within this framework, we establish: (i) the existence of infinitely many entire positive radial solutions for a nonempty set of central values; (ii) closedness of the set of admissible central values; and (iii) largeness (blow-up at infinity) of solutions corresponding to boundary points of this set. The approach is based on a novel subharmonic functional tailored to the reciprocal integral conditions, extending classical Keller–Osserman theory to a broad class of coupled systems with general nonlinearities and weight functions.
MSC 2020: 35J47, 35B40, 35B09, 35J60.
Keywords and phrases: Semilinear elliptic systems, Keller–Osserman condition, radial solutions, large solutions, blow-up at infinity.
@@ 1 Introduction
for which Keller [5] and Osserman [6] established the celebrated Keller–Osserman condition, providing sharp integral criteria for the existence of large solutions (see also [11] fro more details). These foundational results have inspired extensive research on systems of elliptic equations, where the interaction between components introduces new analytical difficulties.
@@
Remark 1 (Relation to Lair’s result).
@@
Let $S$ be the set of all central values $(a,b)\in[0,\infty)^{2}$ for which (1) admits an entire positive radial solution. Then $S$ is nonempty and closed.
@@
Paper structure. Section 2 contains some preliminaries, including the radial formulation of the system, local existence results, and the Keller–Osserman transform, which will be used in the proofs. Section 3 proves the existence of entire positive radial solutions for a nonempty set of central values. Section 4 shows that the set of all central values of entire solutions is nonempty and closed. Section 5 establishes that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity. Section 6 presents concluding remarks and possible extensions. An Appendix contains an auxiliary proposition on the integrability of compositions of the nonlinearities.
@@ 2 Preliminaries: radial formulation, local existence and Keller–Osserman transform
For any $a,b>0$ there exists $\rho=\rho(a,b)>0$ and a positive solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (4) with the given central values. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r\in(0,\rho)$.
@@
Finally, $p,q\geq 0$ and $f,g\geq 0$ imply $u^{\prime},v^{\prime}$ are nondecreasing with $u^{\prime}(0)=v^{\prime}(0)=0$, hence $u^{\prime},v^{\prime}\geq 0$ on $(0,\rho)$. ∎
@@
Let $(u,v)$ be any positive radial $C^{2}$-solution of (4). Then, for every $r>0$, one has
@@
$v(r)\leq b+\mathcal{L}_{q}\,f(u(r)),$
where $\mathcal{L}_{q}$ is the finite constant defined in (3), i.e.
$\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds.$
Since $g$ is nondecreasing and positive, we can estimate
@@
By the multiplicative subadditivity of $g$,
$g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right)\leq g\!\left(f(u(r))\right)\cdot g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right).$
Moreover, since $u(r)\geq a$ for all $r\geq 0$ by monotonicity, we have
$\frac{b}{f(u(r))}+\mathcal{L}_{q}\ \leq\ \frac{b}{f(a)}+\mathcal{L}_{q},$
and thus
@@
A completely analogous argument, starting from
$u(r)\leq a+\mathcal{L}_{p}\,g(v(r)),$
and using the multiplicative subadditivity of $f$, gives
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq\ q(r)\,f(g(v(r)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
which is the second inequality in (9).
We now derive (10). By the chain rule,
$\Phi^{\prime}(u)=-\frac{u^{\prime}}{g(f(u))},$
and differentiating again,
$\Phi^{\prime\prime}(u)=-\frac{u^{\prime\prime}}{g(f(u))}+\frac{g^{\prime}(f(u))\,f^{\prime}(u)}{[g(f(u))]^{2}}\,(u^{\prime})^{2}.$
Since $f^{\prime},g^{\prime}\geq 0$, the second term is nonnegative, hence
$\Phi^{\prime\prime}(u)\ \geq\ -\frac{u^{\prime\prime}}{g(f(u))}.$
Adding $\frac{n-1}{r}\Phi^{\prime}(u)=-\frac{n-1}{r}\frac{u^{\prime}}{g(f(u))}$, we obtain
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -\frac{u^{\prime\prime}+\frac{n-1}{r}u^{\prime}}{g(f(u))}.$
Using the first inequality in (9) and cancelling $g(f(u))$ in numerator and denominator, we deduce
$\Phi^{\prime\prime}(u)+\frac{n-1}{r}\Phi^{\prime}(u)\ \geq\ -p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Similarly, for $\Psi$ we have
$\Psi^{\prime}(v)=-\frac{v^{\prime}}{f(g(v))},\qquad\Psi^{\prime\prime}(v)=-\frac{v^{\prime\prime}}{f(g(v))}+\frac{f^{\prime}(g(v))\,g^{\prime}(v)}{[f(g(v))]^{2}}\,(v^{\prime})^{2}\ \geq\ -\frac{v^{\prime\prime}}{f(g(v))}.$
Adding $\frac{n-1}{r}\Psi^{\prime}(v)$ and using the second inequality in (9) yields
$\Psi^{\prime\prime}(v)+\frac{n-1}{r}\Psi^{\prime}(v)\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$
This completes the proof. ∎
@@ 3 Existence of entire positive solutions
$\mathcal{L}_{p}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}p(\tau)\,d\tau\right)ds,\quad\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds,$
@@
Let $(a,b)\in\mathcal{T}$  be fixed. By Lemma 2.1, there exists $\rho>0$ and a unique local positive radial solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (1) with
@@
$R:=\sup\left\{\rho>0\ :\ (u,v)\ \text{exists and is positive on }[0,\rho]\right\}.$
If $R=\infty$, then $(u,v)$ is already an entire positive radial solution and we are done. Suppose, for contradiction, that $R<\infty$.
@@
$R_{0}:=\sup\left\{\rho\in(0,R]\ :\ u(r)<z_{1}(r)\ \text{and}\ v(r)<z_{2}(r)\ \forall r\in[0,\rho]\right\}.$
@@
Integral formulation and comparison. For $r\in[0,R_{0}]$, the integral form of (1) gives(5)-(6). Since $v(s)<z_{2}(s)$ for $s\leq R_{0}$ and $g$ is increasing, we have
$g(v(s))<g(z_{2}(s)).$
Moreover, by the structure of (11) and the constants $\mathcal{L}_{p},\mathcal{L}_{q}$, we can bound
$g(v(s))\leq g\!\left(f(u(s))\right)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right),$
and similarly for $f(u(s))$. Thus, at $r=R_{0}$,
$u(R_{0})$ $=a+\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\right)dt$
$\leq a+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,ds\right)dt$
$<c+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}p(s)\,g(f(z_{1}(s)))\,ds\right)dt$
$=z_{1}(R_{0}),$
where the strict inequality follows from $u(s)<z_{1}(s)$ for $s<R_{0}$ and the monotonicity of $g\circ f$.
Similarly,
$v(R_{0})$ $=b+\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\right)dt$
$\leq b+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\right)dt$
$<d+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{R_{0}}t^{1-n}\left(\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\right)dt$
$=z_{2}(R_{0}).$
Thus
$u(R_{0})<z_{1}(R_{0}),\quad v(R_{0})<z_{2}(R_{0}).$
By continuity, there exists $\delta>0$ such that the inequalities persist on $0,R_{0}+\delta]$, contradicting the definition of $R_{0}$. Thus $R_{0}=R$ and (12) holds.
Step 2: Extension beyond $R$. From (12) and the fact that $z_{1}$, $z_{2}$ are finite on $[0,R]$, we have $u(R^{-})<\infty$ and $v(R^{-})<\infty$. By monotonicity of $u,v$ (Lemma 2.1), $u,v$ and their derivatives are bounded on $[0,R)$. Standard continuation for ODEs then extends $(u,v)$ beyond $R$, contradicting the definition of $R$. Hence $R=\infty$ and $(u,v)$ is an entire positive radial solution. ∎

### v2

@@ Abstract
We investigate the existence, multiplicity, and asymptotic behavior of entire positive radial solutions to the semilinear elliptic system
@@
under new Keller–Osserman-type integral conditions on the nonlinearities $f$ and $g$, and decay constraints on the radial weight functions $p$ and $q$. The nonlinearities are assumed continuous on $0,\infty)$, differentiable on $(0,\infty)$, nondecreasing, multiplicatively subadditive, vanishing at the origin, and strictly positive elsewhere, satisfying the finite reciprocal integral conditions
[…]
The radial weights satisfy $\int_{0}^{\infty}s\,p(s)\,ds<\infty$ and $\int_{0}^{\infty}s\,q(s)\,ds<\infty$, with $\min(p,q)$ not compactly supported. Within this framework, we establish three main results: (i) existence of infinitely many entire positive radial solutions for each central value $(a,b)$ in a nonempty open set $\mathcal{T}\subset(0,\infty)^{2}$; (ii) closedness of the set $S$ of all admissible central values in $0,\infty)^{2}$; and (iii) largeness (blow-up at infinity) of solutions corresponding to boundary points $E=\partial S\cap(0,\infty)^{2}$. The proofs employ novel comparison techniques with auxiliary scalar problems, Arzelã -Ascoli compactness arguments, and a subharmonic functional approach tailored to the reciprocal integral conditions via Keller–Osserman-type transforms. Our results extend classical Keller–Osserman theory to a broad class of coupled elliptic systems with general nonlinearities and weight functions, unifying and generalizing several existing frameworks.
MSC 2020: 35J47 (Primary), 35B40, 35B09, 35J60, 35B33 (Secondary).
Keywords: Semilinear elliptic systems; Keller–Osserman condition; entire positive radial solutions; large solutions; blow-up at infinity; multiplicative subadditivity; comparison principle; weighted elliptic systems; Arzelã -Ascoli theorem.
@@ 1 Introduction
for which Keller [5] and Osserman [10] established the celebrated Keller–Osserman condition, providing sharp integral criteria for the existence of large solutions (see also [12] for more details). These foundational results have inspired extensive research on systems of elliptic equations, where the interaction between components introduces new analytical difficulties.
@@
Remark 1.1 (Relation to Lair’s result).
@@
Let $S$ be the set of all central values $(a,b)\in(0,\infty)^{2}$ for which (1) admits an entire positive radial solution. Then $S$ is nonempty and closed.
@@
Paper structure. Section 2 contains some preliminaries, including the radial formulation of the system, local existence results, and the Keller–Osserman transform, which will be used in the proofs. Section 3 proves the existence of entire positive radial solutions for a nonempty set of central values. Section 4 shows that the set of all central values of entire solutions is nonempty and closed. Section 5 establishes that solutions corresponding to boundary points of this set are large, i.e., both components blow up at infinity. Compatibility between Theorem 1.4 and Theorem 2.4 of [3] is given in Section 6. Section 7 presents concluding remarks and possible extensions. An Appendix contains an auxiliary proposition on the integrability of compositions of the nonlinearities.
@@ 2 Preliminaries: radial formulation, local existence and Keller–Osserman transform
Before proving our results, we note first that for $a,b\in\left(0,\infty\right)$ any solution of the system of integral equations (6)–(7) valid for all $r\geq 0$ will also be a positive entire solution of (1) but not necessarily (4).
@@
For any $a,b>0$ there exists $\rho=\rho(a,b)>0$ and a positive solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (5) with the given central values. Moreover, $u^{\prime}(r)\geq 0$ and $v^{\prime}(r)\geq 0$ for all $r\in[0,\rho)$.
@@
Finally, $p,q\geq 0$ and $f,g\geq 0$ imply $u^{\prime},v^{\prime}$ are nondecreasing with $u^{\prime}(0)=v^{\prime}(0)=0$, hence $u^{\prime},v^{\prime}\geq 0$ on $0,\rho)$. ∎
@@
Let $a,b>0$ and $(u,v)$ be any positive radial $C^{2}$-solution of (5). Then, for every $r>0$, one has
@@
$v(r)=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt.$
Since $u(s)\geq u(0)=a>0$ for all $s\geq 0$ by monotonicity, and $f$ is nondecreasing, we obtain
$v(r)\leq b+f(u(r))\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,ds\,dt.$
By the decay assumption (3), the double integral converges:
$\mathcal{L}_{q}:=\frac{1}{n-2}\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds=\lim_{r\to\infty}\mathcal{Q}(r)<\infty.$
Therefore, for all $r\geq 0$,
$v(r)\leq b+\mathcal{L}_{q}\,f(u(r)).$ (12)
Since $g$ is nondecreasing and positive, we estimate
@@
Now we apply the multiplicative subadditivity property (F2). Writing
$b+\mathcal{L}_{q}\,f(u(r))=f(u(r))\cdot\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right),$
and noting that $f(u(r))>0$ for $u(r)>0$, we invoke (F2) with $l=f(u(r))$ and $t=\frac{b}{f(u(r))}+\mathcal{L}_{q}$ to obtain
$g\!\left(b+\mathcal{L}_{q}\,f(u(r))\right)=g\!\left(f(u(r))\cdot\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right)\right)\leq g(f(u(r)))\cdot g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right).$
Since $u(r)\geq a$ for all $r\geq 0$ by monotonicity, and $f$ is nondecreasing, we have $f(u(r))\geq f(a)$, hence
$\frac{b}{f(u(r))}\leq\frac{b}{f(a)},$
and thus by monotonicity of $g$,
$g\!\left(\frac{b}{f(u(r))}+\mathcal{L}_{q}\right)\leq g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Combining these estimates, we obtain
@@
$u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)=p(r)g(v(r))\ \leq\ p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
By complete symmetry, we establish the bound for $u$:
$u(r)\leq a+\mathcal{L}_{p}\,g(v(r)),$
where $\mathcal{L}_{p}$ is defined in (3). Applying the multiplicative subadditivity of $f$ and monotonicity of $f$, $v(r)\geq b$ implies
$f(u(r))\leq f(a+\mathcal{L}_{p}g(v(r)))\leq f(g(v(r)))f\left(\frac{a}{g(v(r))}+\mathcal{L}_{p}\right)\leq f(g(v(r)))f\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
yielding the second inequality in (10):
$v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)\ \leq\ q(r)\,f(g(v(r)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right).$
We now derive (11) from (10). Recall that $\Phi(t)=\int_{t}^{\infty}\frac{ds}{g(f(s))}$ is $C^{1}$ on $(0,\infty)$ with
$\Phi^{\prime}(t)=-\frac{1}{g(f(t))}<0.$
Applying the chain rule to $\Phi\circ u$, we obtain
$\frac{d}{dr}\Phi(u(r))=\Phi^{\prime}(u(r))\,u^{\prime}(r)=-\frac{u^{\prime}(r)}{g(f(u(r)))},$
and differentiating a second time,
$\frac{d^{2}}{dr^{2}}\Phi(u(r))$ $=\Phi^{\prime\prime}(u(r))(u^{\prime}(r))^{2}+\Phi^{\prime}(u(r))u^{\prime\prime}(r)$
$=-\frac{u^{\prime\prime}(r)}{g(f(u(r)))}+\frac{d}{dt}\left[-\frac{1}{g(f(t))}\right]_{t=u(r)}(u^{\prime}(r))^{2}$
$=-\frac{u^{\prime\prime}(r)}{g(f(u(r)))}+\frac{g^{\prime}(f(u(r)))\,f^{\prime}(u(r))}{[g(f(u(r)))]^{2}}\,(u^{\prime}(r))^{2}.$
By assumption (F1), $f^{\prime},g^{\prime}\geq 0$ and $u^{\prime}\geq 0$, so the second term is nonnegative. Hence
$\Phi^{\prime\prime}(u(r))\ \geq\ -\frac{u^{\prime\prime}(r)}{g(f(u(r)))}.$
Moreover,
$\frac{n-1}{r}\Phi^{\prime}(u(r))=-\frac{n-1}{r}\,\frac{u^{\prime}(r)}{g(f(u(r)))}.$
Adding these two inequalities gives
$\Phi^{\prime\prime}(u(r))+\frac{n-1}{r}\Phi^{\prime}(u(r))\ \geq\ -\frac{u^{\prime\prime}(r)+\frac{n-1}{r}u^{\prime}(r)}{g(f(u(r)))}.$ (13)
From the first inequality in (10), we have
@@
Substituting this into (13) and noting that $g(f(u(r)))$ cancels in the numerator and denominator, we deduce
$\Phi^{\prime\prime}(u(r))+\frac{n-1}{r}\Phi^{\prime}(u(r))\ \geq\ -\frac{p(r)\,g(f(u(r)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)}{g(f(u(r)))}=-p(r)\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
This is the first inequality in (11).
By complete symmetry, we have for $\Psi(t)=\int_{t}^{\infty}\frac{ds}{f(g(s))}$:
$\Psi^{\prime}(v(r))=-\frac{v^{\prime}(r)}{f(g(v(r)))},$
and
$\Psi^{\prime\prime}(v(r))=-\frac{v^{\prime\prime}(r)}{f(g(v(r)))}+\frac{f^{\prime}(g(v(r)))\,g^{\prime}(v(r))}{[f(g(v(r)))]^{2}}\,(v^{\prime}(r))^{2}\ \geq\ -\frac{v^{\prime\prime}(r)}{f(g(v(r)))}.$
Thus
$\Psi^{\prime\prime}(v(r))+\frac{n-1}{r}\Psi^{\prime}(v(r))\ \geq\ -\frac{v^{\prime\prime}(r)+\frac{n-1}{r}v^{\prime}(r)}{f(g(v(r)))}.$
Using the second inequality in (10) and canceling $f(g(v(r)))$ yields
$\Psi^{\prime\prime}(v(r))+\frac{n-1}{r}\Psi^{\prime}(v(r))\ \geq\ -q(r)\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
which is the second inequality in (11). This completes the proof. ∎
@@ 3 Existence of entire positive solutions
$\mathcal{L}_{p}:=\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}p(\tau)\,d\tau\right)ds,\quad\mathcal{L}_{q}:=\int_{0}^{\infty}s^{1-n}\left(\int_{0}^{s}\tau^{n-1}q(\tau)\,d\tau\right)ds,$
@@
Let $(a,b)\in\mathcal{T}$ be fixed, i.e., $0<a<c$ and $0<b<d$. By Lemma 2.1, there exists $\rho>0$ and a unique local positive radial solution $(u,v)\in C^{2}([0,\rho))^{2}$ of (1) satisfying the integral equations (6)–(7) with initial values
@@
$R:=\sup\left\{\rho>0\ :\ (u,v)\ \text{exists as a positive $C^{2}$ solution on }[0,\rho]\right\}.$
By Lemma 2.1, $R>0$. If $R=\infty$, then $(u,v)$ is already an entire positive radial solution and we are done.
We now prove that $R=\infty$ by contradiction. Suppose that $R<\infty$.
@@
$R_{0}:=\sup\left\{\rho\in(0,R]\ :\ u(r)<z_{1}(r)\ \text{and}\ v(r)<z_{2}(r)\text{,}\ \forall r\in[0,\rho]\right\}.$
@@
Integral formulation and comparison. For $r\in[0,R_{0}]$, we have from the integral formulation (6)–(7):
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt,$
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt.$
By Lemma 2.3, specifically inequality (10), we have for all $s\in[0,R_{0}]$:
$g(v(s))\leq g(f(u(s)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right).$
Therefore, for any $r\in[0,R_{0}]$,
$u(r)$ $=a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(v(s))\,ds\,dt$
$\leq a+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\,ds\,dt.$ (16)
Recall that $z_{1}$ satisfies (see (14)):
$z_{1}(r)=c+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(\frac{b}{f(a)}+\mathcal{L}_{q})\,g(f(z_{1}(s)))\,ds\,dt.$
Since $u(s)<z_{1}(s)$ for all $s\in[0,R_{0})$ and $g\circ f$ is nondecreasing (by (F1)), we have
$g(f(u(s)))<g(f(z_{1}(s)))\quad\text{for all }s\in(0,R_{0}).$
Moreover, $a<c$ by assumption. Therefore, for $r\in(0,R_{0}]$, from (16):
$u(r)$ $\leq a+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(u(s)))\,ds\,dt$
$<c+g\!\left(\frac{b}{f(a)}+\mathcal{L}_{q}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}p(s)\,g(f(z_{1}(s)))\,ds\,dt$
$=z_{1}(r),$
where the strict inequality uses $a<c$ and $g(f(u(s)))<g(f(z_{1}(s)))$ for $s\in(0,r]$. In particular, taking $r=R_{0}$, we obtain
$u(R_{0})<z_{1}(R_{0}).$
By exactly the same reasoning, using Lemma 2.3 to obtain
$f(u(s))\leq f(g(v(s)))\,f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right),$
and noting that $v(s)<z_{2}(s)$ for $s\in[0,R_{0})$ implies $f(g(v(s)))<f(g(z_{2}(s)))$, and $b<d$, we deduce
$v(r)$ $=b+\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(u(s))\,ds\,dt$
$\leq b+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(g(v(s)))\,ds\,dt$
$<d+f\!\left(\frac{a}{g(b)}+\mathcal{L}_{p}\right)\int_{0}^{r}t^{1-n}\int_{0}^{t}s^{n-1}q(s)\,f(g(z_{2}(s)))\,ds\,dt$
$=z_{2}(r).$
In particular, at $r=R_{0}$,
$v(R_{0})<z_{2}(R_{0}).$
Thus we have shown that
$u(R_{0})<z_{1}(R_{0})\quad\text{and}\quad v(R_{0})<z_{2}(R_{0}).$
By the continuity of $u,v,z_{1},z_{2}$, there exists $\delta>0$ such that
$u(r)<z_{1}(r)\quad\text{and}\quad v(r)<z_{2}(r)\quad\text{for all }r\in[0,R_{0}+\delta],$
contradicting the supremum definition of $R_{0}$. Therefore, we must have $R_{0}=R$ and (15) holds for all $r\in[0,R]$.
Step 2: Extension beyond $R$. From (15), we have
$u(r)\leq z_{1}(r)\quad\text{and}\quad v(r)\leq z_{2}(r)\quad\text{for all }r\in[0,R).$
Since $z_{1}$ and $z_{2}$ are entire positive radial solutions (hence finite on $[0,R]$ for any finite $R$), there exist constants $M_{1},M_{2}>0$ such that
$z_{1}(r)\leq M_{1}\quad\text{and}\quad z_{2}(r)\leq M_{2}\quad\text{for all }r\in[0,R].$
Therefore, by comparison,
$u(r)\leq M_{1}\quad\text{and}\quad v(r)\leq M_{2}\quad\text{for all }r\in[0,R).$
Since $u$ and $v$ are nondecreasing (Lemma 2.1), the limits
$\lim_{r\to R^{-}}u(r)=:u(R^{-})\leq M_{1}<\infty,\quad\lim_{r\to R^{-}}v(r)=:v(R^{-})\leq M_{2}<\infty$
exist and are finite.
Moreover, from the radial equations (5),
$u^{\prime}(r)=r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(v(s))\,ds,$
and since $v(s)\leq M_{2}$ for $s\in[0,R)$ and $g$ is continuous, we have $g(v(s))\leq g(M_{2})$ for all $s\in[0,R)$. Thus
$u^{\prime}(r)\leq r^{1-n}\int_{0}^{r}s^{n-1}p(s)g(M_{2})\,ds\leq g(M_{2})\mathcal{P}(R)<\infty,$
where $\mathcal{P}(R)<\infty$ by assumption (PQ). Similarly, $v^{\prime}(r)\leq f(M_{1})\mathcal{Q}(R)<\infty$ for all $r\in[0,R)$.
Therefore, $(u,v)$ and $(u^{\prime},v^{\prime})$ are uniformly bounded on $[0,R)$. By the fundamental theorem of calculus and uniform continuity, the limits
$\lim_{r\to R^{-}}(u(r),u^{\prime}(r),v(r),v^{\prime}(r))$
exist and are finite. By the standard existence and uniqueness theory for ODEs (Lemma 2.1 applied with initial data at $r=R$), the solution can be extended beyond $R$ to some interval $0,R+\varepsilon)$ with $\varepsilon>0$. This contradicts the maximality of $R$.
Hence, we conclude that $R=\infty$, and $(u,v)$ is an entire positive radial solution of (1) with $u(0)=a$, $v(0)=b$, and $u^{\prime}(r)\geq 0$, $v^{\prime}(r)\geq 0$ for all $r\geq 0$. ∎